1.8) a. To estimate the nest success rate of Common Terns in the Lake Ontan pmithologist studies a random sample of 150 nests where 1 proportion of successful broods for this species of bird on Lake 120 of the nests had successful broods d on these results, create a 98 percent confidence interval to estimate the population Ontario 98% Confidence interval To maintain a heathy population of Common Terns, more than 70% of all nests should have successful broods, does the confidence interval you created above seem to indicate a healthy population of Common Terns for the Lake Ontario region? Why or why not? (be brief) b.
Answer #1

Ans;

a)Sample proportion=120/150=0.8

98% confidence interval of population proportion

=0.8+2.326*sqrt(0.8*(1-0.8)/150)

=0.8+/-0.076

=(0.724, 0.876)

or (72.4% , 87.6%)

b)As,lower limit of above confidence interval is higher than 70%,so it indicates that more than 70% of all nests should have successful broods.